use the laplace transfrom method to find y(t) sllution of the initial value problem y''-7y'+12y=0, y(0)=2, y'(0)=-2
Answer #1

Given initial value problem is y- 7y +12y = 0 4102=2 , y(0)=2 Taxe laplace transform on both sieles → 24 g (+1} -7 24 yıtl}Take s=3 -4 -A -> A=4 Take s=4 & -2 B B=-2 4 2 2S-10 al s-3 (5-3)15-21) s-4 (5-3)(3-4) ut 3t -4 e - 2e in e ее. О ut becomes.